A rod of length L is hinged from one end.It is brougt to horizontal position and then released. the angular velocity of the rod when it is in vertical positon

Grade:11

1 Answers

vikas askiitian expert
13 years ago

wrt initial position of rod , when roD is verticle then center of mass of rod has fallen a height of L/2.....

now , applying conservation of energy

  change in (kE) = difference in potential energy

               IW2/2 = mgdh

              (mL2/3)(W2/2) = mg(L/2)                   

             W =(3g/L)1/2

 Irod = mL2/3 , dh = L/2